How many kg of water must be evaporated from 100 kg of a 6% salt solution to obtain a 20% salt solution?
Answer by stanbon(75887)

How many kg of water must be evaporated from 100 kg of a 6% salt solution to obtain a 20% salt solution?
------
Equation:
water - water = water
----
0.94*100 - 1*x = 0.80(100 - x)
-----
94*100 - 100x = 80*100 - 80x
------
-20x = -14*100
x = 70 kg (amt. of water to evaporate)
==============
Cheers,
Stan H.
==============